Rank the following three acids in terms of decreasing acid strength. HBrO4, HBrO2, and HBrO3

a. HBrO3 > HBrO2 > HBrO4

b. HBrO2 > HBrO4 > HBrO3

c. HBrO4 > HBrO3 > HBrO2

d. HBrO3 > HBrO4 > HBrO2

e. HBrO2 > HBrO3 > HBrO4
Answer #1

Gires that HBr Oy, HBrO3; H Broz oxidation state in HBroy of Br HBroy 1+2+(-2) x4=0 . 2= +7 In HBrO3 = 1+2+(-2) x3=0 2-50 x=+Please give better rating